# The Self-Duality Equations on a Riemann Surface and Four-Dimensional Chern–Simons Theory

### v1

@@ 2 From 4d Chern–Simons Theory to Hitchin’s Equations
$\frac{1}{4\pi}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\text{Res}_{z=0}\big[\omega\wedge\mathrm{tr}(\mathsf{A}\wedge\delta\mathsf{A})\big]\,$, (2.13)
and we must impose boundary conditions on the components of $\mathsf{A}$ near $z=0$ so that this vanish for all variations. Thinking of (2.13) as a 1-form on the space of $\mathsf{A}$’s, its vanishing for all variations requires that the germ of $\mathsf{A}$ at $z=0$ lies in a Lagrangian subspace for the skew form defined by its variation
@@
and these are also sufficient to make the action bounded and gauge invariant near the pole at $z=0$. With our $\omega$ as (2.12), we must make such a choice at both $z=0$ and $\infty$. For the reduction to the WZW-model, one takes $p=0$ at $z=0$ and p=1p=-1 at $z=\infty$ [CostelloYamazaki201908]. For the reduction to the Hitchin equations, we impose the boundary conditions
@@
Other natural reductions of the four-dimensional self-duality equations are obtained by different choices of reality condition. For example, straightforward complex conjugation within a complex Lie group rather than hermitian conjugation will give for example $A$ arising as an $\text{so}(n)$ rather than $\text{su}(n)$ gauge connection. Alternatively, one can lift the conjugation $z\rightarrow+1/\bar{z}$ that has fixed points so that on the fixed point set $|z|=1$, $\mathsf{A}$ as a whole then takes values in the real Lie algebra. This gives systemsstart gauge equivalent to harmonic maps. See [MasonWoodhouse199605, §6.2] for related examples. ∎
@@ 2.3 Derivation of the 2d Action from 4d CS Theory
This section contains our main aim, i.e., the derivation of the action (2.8) for Hitchin’s equations (2.1) from the four-dimensional Chern–Simons theory setup described in §2.2. We will adopt the familiar strategy from [CostelloYamazaki201908]: First one enters the gauge in which the $\mathrm{d}\bar{z}$ component of $\mathsf{A}$ vanishes. In this gauge the components of the equations of motion involving $\mathrm{d}\bar{z}$ imply that $\mathsf{A}$ is meromorphic and gauge-equivalent to the Lax form (2.2). This is non-trivial as the Lax form (2.2) does not satisfy (2.16) so the requisite gauge transformation is incompatible with the boundary conditions.
@@
In the following all our considerations will be local on $\Sigma$ and we will abuse notation by using $\Sigma$ for a sufficiently small open subset of $\Sigma$.
@@
Thus $h=h^{*}$ is Hermitian. It follows that, using a $\gamma$-gauge transformation (2.20) on $f$ we have
@@
and this can be used to reduce $h$ to the identity assuming that $h$ is positive definite (otherwise to normal form with $\pm 1$ on the diagonal according to signature). We define this to be a unitary gauge in which the residual $\gamma$-gauge freedom is unitary, $\gamma^{*}\gamma=I$ yielding the residual gauge freedom of the Hitchin equations as normally presented. With this, $\mathsf{A}^{f}$ in turn satisfies $\mathsf{A}^{f}(-1/\bar{z})=-\mathsf{A}^{f*}(z)$ so that $A$ is skew-Hermitian and $\widetilde{\varphi}=\varphi^{*}$ as required.
@@
[…]
@@
are invariant. It further follows that if we fix the $\gamma$-gauge by setting $f_{0}=I$, together with our reality conditions, it follows from (2.28) that $h$ coincides with that introduced in (2.28) and so with our relatity conditions, $h=h^{*}$ and so is Hermitian. However, the gauge freedom $\mathsf{A}\rightarrow g\mathsf{A}g^{-1}-(\mathrm{d}g)g^{-1}$ acts by $f\mapsto fg$ so that
@@
We now consider the reduction of the action of 4d CS theory to the 2d action (2.8) on $\Sigma$ for the Hitchin’s equations. More precisely,
@@
To determine the resulting 2d integrable model we follow the procedure in [CostelloYamazaki201908]. The strategy is to gauge fix to the holomorphic gauge (2.22), and integrate out the $\mathsf{A}^{\prime}$ field, defined in (2.24), to land on (2.8) (otherwise said, $\mathsf{A}^{\prime}$ will be seen to decouple and to vanish on shell). This is complicated by the fact that the gauge transformation $f$ in (2.21) does not satisfy (2.23) so that it generates boundary terms that we now identify. Under the local gauge transformation $f$, with $J$ as in (2.31), the Chern–Simons form (1.2) changes as
@@
We first note that the boundary conditions on $\mathsf{A}^{\prime}$ imply that the contribution from $\omega\wedge\mathrm{d}(JA_{0})$ is smooth at $z=0,\infty$ so is exact and integrates to zero over $\mathbb{CP}^{1}$. We are left with the boundary integral
@@
This is essentially a residue calculation on $\mathbb{CP}^{1}$ with second order poles at $z=0,\infty$ multiplying $\varphi,\widetilde{\varphi}$. This picks out the $z$-derivative of $J$ at these points, together with simple poles multiplying $A$. Noting the opposite signs coming from $z=0,\infty$ and using (2.31) and (2.32), we obtain
@@
In our proof above, gauge invariance is broken by the choice $f_{0}=1$. We can re-instate gauge-invariance by allowing non-trivial $f_{0}$ and using it to transform $A$ to a gauge in which $h$ is constant in standard form for a Hermitian metric when reality conditions are imposed. With this our variables become
$(A,\Psi,\widetilde{\Psi})=(-\bar{\partial}f_{0}f_{0}^{-1}-\partial f_{\infty}f_{\infty}^{-1},f_{\infty}\psi f_{\infty}^{-1},f_{0}\widetilde{\psi}f_{0}^{-1})\,.$ (2.48)
In this gauge we have
@@
[…]
where the brackets denote the inner product with respect to now constant $h$.
In a general unitary gauge, $f_{0}$ must be nontrivial subject to (2.28) with $h=I$. The framework admits $\gamma$-gauge transformations but now with $\gamma$ unitary but otherwise arbitrary. With these choices, only the combination $S_{1}+S_{2}$ is gauge invariant.
@@ 2.4 Symplectic Structure, and Hitchin Hamiltonians
We have derived an action (2.8) whose equations of motion are the Hitchin’s equations. On the other hand, $\mathscr{M}_{\text{H}}(\Sigma,G)$, the space of solutions of Hitchin’s equations, is a hyperkähler manifold which has three distinguished complex structures $\mathcal{I},\mathcal{J},$ and $\mathcal{K}$. Therefore, one may wonder how to derive these symplectic structures from the perspective of 4d Chern–Simons theory. We first show that the symplectic structure derived from 4d Chern–Simons theory with $\omega$ as in (2.12) is the canonical symplectic structure associated with the complex structure $\mathcal{I}$, which is
@@
where $|\mathrm{d}^{2}w|$ denotes the measure corresponding to the type-$(1,1)$ form $\mathrm{i}\,\mathrm{d}w\wedge\mathrm{d}\bar{w}$. [^13] Our convention for the canonical symplectic structures on $\mathscr{M}_{\text{H}}(\Sigma,G)$, given in (2.51) and (2.58), are the same as [KapustinWitten200604, eq. (4.8)]. The difference is that $\phi$ is Hermitian here while it is skew-Hermitian in [KapustinWitten200604]. Therefore, there are some sign differences. For example, the relative sign of (2.51) is plus (++) while it is minus (-) in [KapustinWitten200604, eq. (4.8)]. These signs ensure the reality of the corresponding symplectic structures. Furthermore, we explain how to construct Hitchin Hamiltonians in terms of four-dimensional gauge field.
@@
where we have used $\phi=\varphi+\varphi^{*}$ in the second line. We can immediately recognize that $\Omega_{\text{CS}_{4}}$ coincides with $\Omega_{\mathcal{I}}$, given in (2.51). Thus the space of solutions of the field equations of the action (2.8) leads to the symplectic structure that is the Kahler form for $\mathscr{M}_{\text{H}}(\Sigma,G)$ as a complex manifold in the complex structure $\mathcal{I}$. This is the complex structure that arises from the identification of $\mathscr{M}_{\text{H}}(\Sigma,G)$ with the the moduli space of Higgs bundles. We now elaborate further on this.
@@
In [Hitchin198701], Hitchin shows that for compact $\Sigma$ of genus $g\geq 2$, the space of solutions of the action (2.8) is essentially the moduli space $\mathscr{M}_{\text{Hig}}^{\text{s}}(n,d)$ of Higgs bundles, i.e., pairs $(E,\varphi)$ where $E$ is a holmorpihc vector bundle and $\varphi$ the holomorphic $(1,0)$-form with values in $\mathrm{End}(E)$ introduced above. He goes on to show that this moduli space can be realized as an algebraic completely integrable system. We sketch the construction very briefly as follows, and refer to the original paper for the many details concerning stability, completeness etc..
Firstly general hyperkähler lore shows that the complex 2-form form $\Omega_{\mathcal{I}}+\mathrm{i}\Omega_{\mathcal{J}}$ is holomorphic symplectic with respect to the complex structure $\mathcal{I}$. This can be represented explicitly as
$\Omega_{\mathcal{J}}+\mathrm{i}\,\Omega_{\mathcal{K}},=-\frac{1}{2\pi\mathrm{i}}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits(\delta\phi\wedge(1-\mathrm{i}\star)\delta A)=\frac{1}{\pi}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\,|\mathrm{d}^{2}w|\mathrm{tr}(\delta\phi_{w}\wedge\delta A_{\bar{w}}).$ (2.57)
@@
We must then present the complete family of holomorphic commuting Hamiltonians to realize this as an algebraic completely integrable system. These can be presented via the Hitchin fibration
$\mathbf{h}:\mathscr{M}_{\text{Hig}}^{\text{s}}(n)\to\mathbf{B},$ (2.59)
@@
The integrand is now a $(1,1)$-form so it can be integrated over $\Sigma$. The number of independent Hamiltonians corresponds to the number of independent $v$s, which can be computed by the Riemann–Roch Theorem. This gives for $g\geq 2$
@@
Thus we have constructed the Hitchin Hamiltonians from four-dimensional Chern–Simons theory.
3 $\mathbb{CP}^{1}$-Family of 2d Actions Associated with the Hitchin’s Equations
In §2, we described an action whose field equations are the Hitchin’s equations, and connected the space of solutions of its field equations with the moduli space of Higgs bundles. On the other hand, $\mathscr{M}_{\text{H}}(\Sigma,G)$, as a hyperkähler manifold, has a whole $\mathbb{CP}^{1}$-family of symplectic structures (1.4). Therefore, a natural follow-up question is whether there exists a corresponding family of 2d actions such that the space of solutions of their field equations is related to $\mathscr{M}_{\text{H}}(\Sigma,G)$ as a complex manifold in the complex structure $\mathcal{J}_{\zeta}$, given in (1.4).

### v2

@@ 2 From 4d Chern–Simons Theory to Hitchin’s Equations
$\frac{1}{4\pi}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\text{Res}_{z=0}\big[\mathrm{tr}(\mathsf{A}\wedge\delta\mathsf{A})\big]\,$, (2.13)
and we must impose boundary conditions on the components of $\mathsf{A}$ near $z=0$ so that this vanishes for all variations. Thinking of (2.13) as a 1-form on the space of $\mathsf{A}$’s, its vanishing for all variations requires that the germ of $\mathsf{A}$ at $z=0$ lies in a Lagrangian subspace for the skew form defined by its variation
@@
and these are also sufficient to make the action bounded and gauge invariant near the pole at $z=0$. With our $\omega$ as (2.12), we must make such a choice at both $z=0$ and $\infty$. For the reduction to the WZW model, one takes $p=0$ at $z=0$ and p=1p=-1 at $z=\infty$ [CostelloYamazaki201908]. For the reduction to the action for Hitchin’s equations, we impose the boundary conditions
@@
Other natural reductions of the four-dimensional self-duality equations are obtained by different choices of reality condition. For example, straightforward complex conjugation within a complex Lie group rather than hermitian conjugation will give, for example, $A$ arising as an $\text{so}(n)$ rather than $\text{su}(n)$ gauge connection. Alternatively, one can lift the conjugation $z\rightarrow+1/\bar{z}$ that has fixed points so that on the fixed point set $|z|=1$, $\mathsf{A}$ as a whole then takes values in the real Lie algebra. This gives systems that are gauge equivalent to harmonic maps. See [MasonWoodhouse199605, §6.2] for related examples. ∎
@@ 2.3 Derivation of the 2d Action from 4d CS Theory
This section contains our main aim, i.e., the derivation of the action (2.8) for Hitchin’s equations (2.1) from the four-dimensional Chern–Simons theory setup described in §2.2. We will adopt the familiar strategy from [CostelloYamazaki201908]: First, one enters the gauge in which the $\mathrm{d}\bar{z}$ component of $\mathsf{A}$ vanishes. In this gauge the components of the equations of motion involving $\mathrm{d}\bar{z}$ imply that $\mathsf{A}$ is meromorphic and gauge-equivalent to the Lax form (2.2). This is non-trivial as the Lax form (2.2) does not satisfy (2.16), so the requisite gauge transformation is incompatible with the boundary conditions.
@@
In the following, all our considerations will be local on $\Sigma$ and we will abuse notation by using $\Sigma$ for a sufficiently small open subset of $\Sigma$.
@@
Thus $h=h^{*}$, and hence is Hermitian. It follows that, using a $\gamma$-gauge transformation (2.20) on $f$ we have
@@
and this can be used to reduce $h$ to the identity, assuming that $h$ is positive definite (otherwise to normal form with $\pm 1$ on the diagonal according to signature). We define this to be a unitary gauge in which the residual $\gamma$-gauge freedom is unitary, $\gamma^{*}\gamma=I$, yielding the residual gauge freedom of the Hitchin equations as normally presented. With this, $\mathsf{A}^{f}$ in turn satisfies $\mathsf{A}^{f}(-1/\bar{z})=-\mathsf{A}^{f*}(z)$ so that $A$ is skew-Hermitian and $\widetilde{\varphi}=\varphi^{*}$ as required.
@@
Thus $(\partial_{\bar{w}}\widetilde{\psi},\partial_{w}\psi)$ are covariant under (LABEL:psi-gauge).
@@
are invariant. It further follows that if we fix the $\gamma$-gauge by setting $f_{0}=I$, together with our reality conditions, it follows from (2.28) that $h$ coincides with that introduced in (2.28) and so with our reality conditions, $h=h^{*}$ and so is Hermitian. However, the gauge freedom $\mathsf{A}\rightarrow g\mathsf{A}g^{-1}-(\mathrm{d}g)g^{-1}$ acts by $f\mapsto fg$ so that
@@
then we see that $J^{\prime}=J+\mathcal{A}$ with $\mathcal{A}$ as defined in (2.26) cancelling the deviation of $J$ from (2.16).
@@
We now consider the reduction of the action of 4d CS theory to the 2d action (2.8) on $\Sigma$ for Hitchin’s equations. More precisely,
@@
To determine the resulting 2d integrable model, we follow the procedure in [CostelloYamazaki201908]. The strategy is to gauge fix to the holomorphic gauge (2.22), and integrate out the $\mathsf{A}^{\prime}$ field, defined in (2.24), to land on (2.8) (otherwise said, $\mathsf{A}^{\prime}$ will be seen to decouple and to vanish on shell). This is complicated by the fact that the gauge transformation $f$ in (2.21) does not satisfy (2.23) so that it generates boundary terms that we now identify. Under the local gauge transformation $f$, with $J$ as in (2.31), the Chern–Simons form (1.2) changes as
@@
We first note that the boundary conditions on $\mathsf{A}^{\prime}$ imply that the contribution from $\omega\wedge\mathrm{d}(JA_{0})$ is smooth at $z=0,\infty$, so is exact and integrates to zero over $\mathbb{CP}^{1}$. We are left with the boundary integral
@@
This is essentially a residue calculation on $\mathbb{CP}^{1}$ with second-order poles at $z=0,\infty$ multiplying $\varphi,\widetilde{\varphi}$. This picks out the $z$-derivative of $J$ at these points, together with simple poles multiplying $A$. Noting the opposite signs coming from $z=0,\infty$ and using (2.31) and (2.32), we obtain
@@
In our proof above, gauge invariance is only broken when we make the choice $f_{0}=I$. We can relax this condition, allowing a non-trivial $f_{0}$, and when reality conditions are imposed, a natural restriction is one in which $h$ is taken to be in a constant standard form for the Hermitian metric, i.e., the identity matrix when positive definite. The gauge connection $A$ is then skew-Hermitian. With this choice, our variables become
$(A,\Psi,\widetilde{\Psi})=(-\bar{\partial}f_{0}f_{0}^{-1}-\partial f_{\infty}f_{\infty}^{-1},f_{\infty}\psi f_{\infty}^{-1},f_{0}\widetilde{\psi}f_{0}^{-1})\,$, (2.48)
where the Higgs field is
@@
The action (1.6) can be expressed in terms of the gauge-covariant data $f_{0},f_{\infty},\Psi,\widetilde{\Psi}$ by substituting $h=f_{\infty}^{-1}f_{0}$, $\psi=f_{\infty}^{-1}\Psi f_{\infty}$ and $\widetilde{\psi}=f_{0}^{-1}\widetilde{\Psi}f_{0}$, yielding
$S_{\mathrm{H}}[f_{0},f_{\infty},\Psi,\widetilde{\Psi}]=\mathrm{i}\int_{\Sigma}\mathrm{tr}(j_{0}^{1,0}\wedge j_{0}^{0,1}-2j_{\infty}^{1,0}\wedge j_{0}^{0,1}+j_{\infty}^{1,0}\wedge j_{\infty}^{0,1})+I_{\mathrm{WZ}}[f_{\infty},f_{0}]+\int_{\Sigma}|\partial_{A}\Psi|^{2}\,.$ (2.50)
Here $j_{0}=\mathrm{d}f_{0}f_{0}^{-1},j_{\infty}=\mathrm{d}f_{\infty}f_{\infty}^{-1}$. The norm $|\cdot|$ is now defined using the standard constant Hermitian metric, and the Wess–Zumino term takes the same form as in equation (2.10) though now defined using a homotopy $f:\Sigma\times[0,1]$ from $f_{0}$ on $\Sigma\times 0$ to $f_{\infty}$ on $\Sigma\times 1$.
In a general unitary gauge, $f_{0}$ must be nontrivial with $f_{\infty}^{*}f_{0}=I$. These reality conditions are compatible with unitary $\gamma$-gauge transformations, under which the action (LABEL:eq:unitary-action) is invariant.
@@ 2.4 Symplectic Structure, and Hitchin Hamiltonians
We have derived an action (2.8) whose equations of motion are Hitchin’s equations. On the other hand, $\mathscr{M}_{\text{H}}(\Sigma,G)$, the space of solutions of Hitchin’s equations, is a hyperkähler manifold which has three distinguished complex structures $\mathcal{I},\mathcal{J},$ and $\mathcal{K}$. Therefore, one may wonder how to derive these symplectic structures from the perspective of 4d Chern–Simons theory. We first show that the symplectic structure derived from 4d Chern–Simons theory with $\omega$ as in (2.12) is the canonical symplectic structure associated with the complex structure $\mathcal{I}$, and is given by
@@
where $|\mathrm{d}^{2}w|$ denotes the measure corresponding to the type-$(1,1)$ form $\mathrm{i}\,\mathrm{d}w\wedge\mathrm{d}\bar{w}$. [^13] Our convention for the canonical symplectic structures on $\mathscr{M}_{\text{H}}(\Sigma,G)$, given in (2.51) and (2.58), are the same as [KapustinWitten200604, eq. (4.8)]. The difference is that $\phi$ is Hermitian here while it is skew-Hermitian in [KapustinWitten200604]. Therefore, there are some sign differences. For example, the relative sign of (2.51) is plus (++) while it is minus (-) in [KapustinWitten200604, eq. (4.8)]. These signs ensure the reality of the corresponding symplectic structures. Furthermore, we explain how to construct Hitchin Hamiltonians in terms of four-dimensional gauge fields.
@@
where we have used $\phi=\varphi+\varphi^{*}$ in the second line. We immediately recognize that $\Omega_{\text{CS}_{4}}$ coincides with $\Omega_{\mathcal{I}}$, given in (2.51). Thus the space of solutions of the field equations of the action (2.8) leads to the symplectic structure that is the Kähler form for $\mathscr{M}_{\text{H}}(\Sigma,G)$ as a complex manifold in the complex structure $\mathcal{I}$. This is the complex structure that arises from the identification of $\mathscr{M}_{\text{H}}(\Sigma,G)$ with the moduli space of Higgs bundles. We now elaborate further on this.
@@
In [Hitchin198701], Hitchin shows that for compact $\Sigma$ of genus $g\geq 2$, the space of solutions of the action (2.8) is essentially the moduli space $\mathscr{M}_{\text{Hig}}^{\text{s}}(n,d)$ of Higgs bundles, i.e., pairs $(E,\varphi)$ where $E$ is a holomorphic vector bundle and $\varphi$ the holomorphic $(1,0)$-form with values in $\mathrm{End}(E)$ introduced above. He goes on to show that this moduli space can be realized as an algebraically completely integrable Hamiltonian system. We sketch the construction very briefly as follows. The details can be found in [Hitchin198701].
Firstly general hyperkähler lore shows that the complex 2-form $\Omega_{\mathcal{I}}+\mathrm{i}\Omega_{\mathcal{J}}$ is holomorphic symplectic with respect to the complex structure $\mathcal{I}$. This can be represented explicitly as
$\Omega_{\mathcal{J}}+\mathrm{i}\,\Omega_{\mathcal{K}}=-\frac{1}{2\pi\mathrm{i}}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits(\delta\phi\wedge(1-\mathrm{i}\star)\delta A)=\frac{1}{\pi}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\,|\mathrm{d}^{2}w|\mathrm{tr}(\delta\phi_{w}\wedge\delta A_{\bar{w}}).$ (2.57)
@@
We must then present the complete family of holomorphic commuting Hamiltonians to realize this as an algebraically completely integrable system. These can be presented via the Hitchin fibration
$\mathbf{h}:\mathscr{M}_{\text{Hig}}^{\text{s}}(n,d)\to\mathbf{B},$ (2.59)
@@
The integrand is now a $(1,1)$-form, so it can be integrated over $\Sigma$. The number of independent Hamiltonians corresponds to the number of independent $v$s, which can be computed by the Riemann–Roch Theorem. This gives for $g\geq 2$
@@
Thus, we have constructed the Hitchin Hamiltonians from four-dimensional Chern–Simons theory.
3 $\mathbb{CP}^{1}$-Family of 2d Actions Associated with Hitchin’s equations
In §2, we described an action whose field equations are Hitchin’s equations, and connected the space of solutions of its field equations with the moduli space of Higgs bundles. On the other hand, $\mathscr{M}_{\text{H}}(\Sigma,G)$, as a hyperkähler manifold, has a whole $\mathbb{CP}^{1}$-family of symplectic structures (1.4). Therefore, a natural follow-up question is whether there exists a corresponding family of 2d actions such that the space of solutions of their field equations is related to $\mathscr{M}_{\text{H}}(\Sigma,G)$ as a complex manifold in the complex structure $\mathcal{J}_{\zeta}$, given in (1.4).